OpenStudy (anonymous):

At your favorite taco place you can choose flour or corn tortilla mild salsa medium salsa spicy salsa or pico de gallo cheddar or Monterey Jack cheese. How many different tacos can you order

OpenStudy (anonymous):

just a simple answer would be nice please

OpenStudy (kropot72):

With each of the 2 choices of tortilla there are 3 choices of salsa giving 2 * 3 possibilities. And with each of these 2 * 3 possibilities there are 2 choices of cheese giving a total of 2 * 3 * 2 different tacos.

OpenStudy (anonymous):

you forgot the answer dude

OpenStudy (anonymous):

in choosing a vacation spot you narrow down your choice to 2 states each of which has 12 counties and each county has 15 cities. How many different cities are possible you to visit

OpenStudy (kropot72):

Can you work out this product to find the answer? \[2\times 3\times 2=?\]

OpenStudy (shane_b):

If you read @kropot72's first response carefully, you'll see that the answers aren't very complicated and you should be able to see how to do these very quickly on your own.

OpenStudy (anonymous):

true hahahahhahaha except i some how got the wrong answer hahahah and it was 16

OpenStudy (kropot72):

Each state has 12 counties and each of these counties has 15 cities. Therefore each state has 12 * 15 cities. There are 2 states each with 12 * 15 cities, therefore the total number of different cities to visit is \[2\times12\times 15=?\]